Graham is the beneficiary of an annuity due. At an annual effective interest rate of 5%, the present value of payments is 123,000. Tyler uses the first-order Macaulay approximation to estimate the present value of Graham's annuity due at an annual effective interest rate 5.4%. Tyler estimates the present value to be 121,212. Calculate the modified duration of Graham's annuity at 5%.
Answer #1

PV of payments at 5.4% = PV of payments at 5% * [( 1+0.05) / ( 1+ 0.054) ^ Macaulay Duration]

121,212 = 123,000 * [(1.05 / 1.054) ^ Macaulay Duration]

Macaulay Duration = ln (121,212 / 123,000) / ln (1.05 / 1.054)

= 3.8512

Modified Duration = Macaulay Duration / (1+YTM)

Therefore Modified Duration = 3.8512 / 1.05 = 3.67